# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ Appendix E Uniqueness results
In this section we prove a number of uniqueness results. Here we will allow translation to permute anyons, so it may no longer be the case that $a=\mathcal{T}a$.
In the proofs we will need the following fact:
@@
Here we also introduce the background gauge field $A$. Gauging the $\mathbb{Z}_{N}$ symmetry, we add to the Lagrangian the BF term $\frac{N}{2\pi}AdB$. We use equation of motion to integrate out $a_{1}$ and set $a_{2}=-A$, after which the theory becomes
@@
which is precisely the K matrix given in Eq. (63).
@@ E.1 The case of $q$ odd
Let $\mathcal{A}$ be a fermionic Abelian TO of size $|\mathcal{A}|=2q^{2}$ that has an order $q$ boson, where $q$ is odd. If $\mathcal{A}$ is further time reversal invariant then it must be $\mathbb{Z}_{q}$ gauge theory fused with the fundamental fermion.
@@
Here we have set the diagonal element to be an even, which is always possible for $q$ odd. We will show that the only $K$ matrix of this form consistent with time reversal is the one where $n=0\pmod{q}$.
@@
Now we will use the fact that $\mathcal{T}$ is an order two operation. This means that:
@@
We therefore conclude that for $q$ odd the only fermionic Abelian TO, $\mathcal{A}$, which is time reversal invariant, has an order $q$ boson, and is of size $|\mathcal{A}|=2q^{2}$ is given by the $K$-matrix
@@
Let $\mathcal{A}\in mG_{\rm GSD}$ for the group $G$ in Eq. (4), where $q$ is odd. Then $\mathcal{A}$ must be $\mathbb{Z}_{q}$ gauge theory stacked with $\{1,c\}$.
@@
We showed in the main text that there must exist a bosonic vison $v\in\mathcal{A}$ which has order $q$. We further showed that $|\mathcal{A}|=2q^{2}$ and was Abelian if $\mathcal{A}\in mG_{\rm GSD}$. Applying the results of the previous theorem gives us our proof. ∎
@@ E.2 The case of $q$ even
We first want to prove that there exists an order $2q$ boson in this theory, $\gamma$, even if translation permutes anyons. In fact we will be able to do better and find a unique such boson. Since our fermionic TO is Abelian we can decompose $\mathcal{A}=\mathcal{A}_{b}\boxtimes\{1,c\}$, where $\mathcal{A}_{b}$ is a bosonic TO which is closed under fusion. In App. C.3 we reviewed the proof that there is a unique $\gamma\in\mathcal{A}_{b}$ which has the property that:
@@
Let $\mathcal{A}\in mG_{GSD}$ for the group $G$ in Eq. (4). Then $\mathcal{A}$ is made up entirely of: powers of $a$, the unique $\gamma$ discussed above, and $c$, i.e.
@@
First recall from Thm. C.27 in App. C that $\mathcal{A}=\mathcal{A}_{b}\boxtimes\{1,c\}$, where $\mathcal{A}_{b}$ is closed under fusion. We already know that $\gamma\in\mathcal{A}_{b}$. Since $q$ is even it must always be the case that $a^{q}\in\mathcal{A}_{b}$. Suppose then that $a\notin\mathcal{A}_{b}$, then $ac\in\mathcal{A}_{b}$. Since $\mathcal{A}_{b}$ is closed under fusion $a^{q+1}c\in\mathcal{A}_{b}$. We see that $a^{q+1}c$ is an order $2q$ anyon with charge exactly $1/q$ modulo two. Then without loss of generality we can assume $a\in\mathcal{A}_{b}$, since all we know about $a$ are these two facts.
@@
Let $\mathcal{A}\in mG_{\rm GSD}$ for the group $G$ in Eq. (4). Then if the unique $\gamma$ discussed earlier is not a boson $a^{q}$ must be a fermion. This further implies that: $2q\theta_{a}=\pi$, $q/2$ must be odd, and $\theta_{a^{q+1}}=\theta_{a}$.
@@
To show this we will again use the relation
@@
proven in Lapa and Levin (2019) for any fermionic TO. Time reversal invariance will demand that the left hand side is equal to one. Then it is a simple matter of evaluating the right hand side for all anyons; these are of the form $a^{r}\gamma^{s}c^{n}$ by Lem. 79.
For compactness we again write $\tilde{\theta}_{b}=\theta_{b}+\pi Q_{b}$. Then since $\pi Q_{\gamma}=\theta_{\gamma,\gamma}=2\theta_{\gamma}$ we will have that:
@@
If we then sum over $s^{\prime}=0,\ldots,q-1$ we will see that $r=0,q$ are the only options that do not evaluate to zero. The relation will then become:
@@
Let $\mathcal{A}\in mG_{\rm GSD}$ for the group $G$ in Eq. (4). Then the unique $\gamma$ discussed earlier must be a boson or fermion and there is thus an order $2q$ boson in the theory.
@@
We now want to establish the possibilities that $a$ can be mapped to under time reversal. By Lem. 79 we know that all anyons in $\mathcal{A}$ can be written as:
@@
It must also be the case that $-\theta_{a}=\theta_{\mathcal{T}a}$. We will now show that none of the possibilities mentioned above can satisfy this property.
@@
Let $\mathcal{A}$ be a fermionic Abelian TO of size $|\mathcal{A}|=8q^{2}$ that has an order $2q$ boson. If $\mathcal{A}$ is time reversal invariant then it must either be $\mathbb{Z}_{2q}$ gauge theory or $\mathrm{U}(1)_{2q}\boxtimes\mathrm{U}(1)_{-2q}$ (stacked with $\{1,c\}$).
@@
The anyons $d\in\mathcal{A}$ in this theory can again be written as $d=(l,m,k)$ for some integers $l,m,k$ where $k=0,1$. They will have topological spin given by:
@@
Similarly if we have that $\mathcal{T}(0,1,0)=(l^{\prime},m^{\prime},k^{\prime})$ then we will have that:
@@
Suppose that $n=q/2,3q/2,5q/2,7q/2\pmod{4q}$ then the requirement that $n(l^{2}+1)=0\pmod{4q}$ will mean that $l^{2}+1=0\pmod{8}$. Similarly if $n=q,3q$, then we will need $l^{2}+1=0\pmod{4}$. Both obviously imply $l$ must be odd, i.e. $l=2l^{\prime}+1$ for some integer $l^{\prime}$. But then it can easily be checked that $l^{2}+1=2\pmod{8}$ and thus $l^{2}+1\neq 0$ either modulo $4$ or $8$. So then we must have that $n=0,2q$. Clearly $n=0$ is simply $\mathbb{Z}_{2q}$ gauge theory, which is consistent with time reversal.
Suppose now that $n=2q$. Then our $K$-matrix is given by:
@@
$W=\begin{pmatrix}1&0&0\\ 1&1&0\\ 0&0&1\end{pmatrix},\ W\in{\rm GL}(2,\mathbb{Z}),$ (94)
@@
Let $\mathcal{A}\in mG_{GSD}$ for the group $G$ in Eq. (4). Then it must be the case that $a^{2}$ is invariant under translation, i.e. $T_{x}a^{2}=a^{2}=T_{y}a^{2}$. This then implies that $a^{2}$ is a boson. It will also further imply that the unique $\gamma$ discussed earlier is a boson.
@@
since $02\theta_{\gamma}=$ 0. Now we note that $a^{2q}=1$ so it must braid trivially with $a$, i.e. $0=\theta_{a,a^{2q}}=4q\theta_{a}$. Since $q$ is even this means that $(2q^{2}+4q)\theta_{a}=0$. Thus we have that $0=2\pi m/q$ so $m$ must be a multiple of $q$. The only anyons that $a$ can be taken to under translation are thus:
@@
In any case we see that $T_{x}a^{2}=(T_{x}a)(T_{x}a)=a^{2}$. The same logic will hold for $T_{y}$. Thus we see that $a^{2}$ must be invariant under translation  [^18] Indeed, in this theory the same logic will show that $a^{2}$ must be invariant under any unitary symmetry..
Next we will appeal to the logic of App. D.1. There we demonstrated that
@@
Let $\mathcal{A}\in mG_{\rm GSD}$ for the group $G$ in Eq. (4), where $q$ is even. Then $\mathcal{A}$ must be $\mathbb{Z}_{2q}$ gauge theory stacked with the fundamental fermion.

### v2

@@ Appendix E Uniqueness results
In this section we prove a number of uniqueness results. Here we allow translation to permute anyons, so it may no longer be the case that $a=\mathcal{T}a$.
In the proofs we need the following fact:
@@
Here we also introduce the background gauge field $A$. Gauging the $\mathbb{Z}_{N}$ symmetry, we add to the Lagrangian the term $\frac{N}{2\pi}AdB$. We use equation of motion to integrate out $a_{1}$ and set $a_{2}=-A$, after which the theory becomes
@@
which is precisely the upper $2\times 2$ block of the $K$ matrix given in Eq. (63). The lower $1$ on the diagonal is due to the stacking of this bosonic theory with $\{1,c\}$.
@@ E.1 The case of $q$ odd
Let $\mathcal{A}$ be a fermionic Abelian TO of size $|\mathcal{A}|=2q^{2}$ that has an order $q$ boson, where $q$ is odd. If $\mathcal{A}$ is further time-reversal invariant then it must be $\mathbb{Z}_{q}$ gauge theory fused with the fundamental fermion.
@@
Here we have set the diagonal element to be an even, which is always possible for $q$ odd. We show that the only $K$ matrix of this form consistent with time reversal is the one where $n=0\pmod{q}$.
@@
Now we use the fact that $\mathcal{T}$ is an order two operation. This means that:
@@
We therefore conclude that for $q$ odd the only fermionic Abelian TO, $\mathcal{A}$, which is time-reversal invariant, has an order $q$ boson, and is of size $|\mathcal{A}|=2q^{2}$ is given by the $K$-matrix
@@
Let $\mathcal{A}\in mG^{(\nu)}_{\rm GSD}$ for the group $G$ in Eq. (4), where $q$ is odd. Then $\mathcal{A}$ must be $\mathbb{Z}_{q}$ gauge theory stacked with $\{1,c\}$.
@@
We showed in the main text that there must exist a bosonic vison $v\in\mathcal{A}$ which has order $q$. We further showed that $|\mathcal{A}|=2q^{2}$ and was Abelian if $\mathcal{A}\in mG^{(\nu)}_{\rm GSD}$. Applying the results of the previous theorem gives us our proof. ∎
@@ E.2 The case of $q$ even
We first want to prove that there exists an order $2q$ boson in this theory, $\gamma$, even if translation permutes anyons. In fact we will be able to do better and find a unique such boson. Since our fermionic TO is Abelian we can decompose $\mathcal{A}=\mathcal{A}_{b}\boxtimes\{1,c\}$, where $\mathcal{A}_{b}$ is a bosonic TO which is closed under fusion. In Appendix C.3 we reviewed the proof that there is a unique $\gamma\in\mathcal{A}_{b}$ which has the property that: [^23] All results thus far about $\gamma$ in Abelian orders are proven in Ref. Lapa and Levin (2019).
@@
Let $\mathcal{A}\in mG^{(\nu)}_{GSD}$ for the group $G$ in Eq. (4). Then $\mathcal{A}$ is made up entirely of: powers of $a$, the unique $\gamma$ discussed above, and $c$, i.e.
@@
First recall from Thm. C.27 in Appendix C that $\mathcal{A}=\mathcal{A}_{b}\boxtimes\{1,c\}$, where $\mathcal{A}_{b}$ is closed under fusion. We already know that $\gamma\in\mathcal{A}_{b}$. Since $q$ is even it must always be the case that $a^{q}\in\mathcal{A}_{b}$. Suppose then that $a\notin\mathcal{A}_{b}$, then $ac\in\mathcal{A}_{b}$. Since $\mathcal{A}_{b}$ is closed under fusion $a^{q+1}c\in\mathcal{A}_{b}$. We see that $a^{q+1}c$ is an order $2q$ anyon with charge exactly $1/q$ modulo two. Then without loss of generality we can assume $a\in\mathcal{A}_{b}$, since all we know about $a$ are these two facts.
@@
Let $\mathcal{A}\in mG^{(\nu)}_{\rm GSD}$ for the group $G$ in Eq. (4). Then if the unique $\gamma$ discussed earlier is not a boson $a^{q}$ must be a fermion. This further implies that: $2q\theta_{a}=\pi$, $q/2$ must be odd, and $\theta_{a^{q+1}}=\theta_{a}$.
@@
To show this we again use the relation
@@
proven in Ref. Lapa and Levin (2019) for any fermionic TO. Time-reversal invariance will demand that the left-hand side is equal to one. Then it is a simple matter of evaluating the right-hand side for all anyons; these are of the form $a^{r}\gamma^{s}c^{n}$ by Lem. 79.
For compactness we again write $\tilde{\theta}_{b}=\theta_{b}+\pi Q_{b}$. Then since $\pi Q_{\gamma}=\theta_{\gamma,\gamma}=2\theta_{\gamma}$ we have that:
@@
If we then sum over $s^{\prime}=0,\ldots,q-1$ we see that $r=0,q$ are the only options that do not evaluate to zero. The relation will then become:
@@
Let $\mathcal{A}\in mG^{(\nu)}_{\rm GSD}$ for the group $G$ in Eq. (4). Then the unique $\gamma$ discussed earlier must be a boson or fermion and there is thus an order $2q$ boson in the theory.
@@
We now want to establish the possibilities that $a$ can be mapped to under time reversal. By Lem. 79 we know that all anyons in $\mathcal{A}$ can be written as
@@
It must also be the case that $-\theta_{a}=\theta_{\mathcal{T}a}$. We now show that none of the possibilities mentioned above can satisfy this property.
@@
Let $\mathcal{A}$ be a fermionic Abelian TO of size $|\mathcal{A}|=8q^{2}$ that has an order $2q$ boson. If $\mathcal{A}$ is time-reversal invariant then it must either be $\mathbb{Z}_{2q}$ gauge theory or $\mathrm{U}(1)_{2q}\boxtimes\mathrm{U}(1)_{-2q}$ (stacked with $\{1,c\}$).
@@
The anyons $d\in\mathcal{A}$ in this theory can again be written as $d=(l,m,k)$ for some integers $l,m,k$ where $k=0,1$. They will have topological spin given by
@@
Similarly if we have that $\mathcal{T}(0,1,0)=(l^{\prime},m^{\prime},k^{\prime})$ then we have that:
@@
Suppose that $n=q/2,3q/2,5q/2,7q/2\pmod{4q}$ then the requirement that $n(l^{2}+1)=0\pmod{4q}$ will mean that $l^{2}+1=0\pmod{8}$. Similarly if $n=q,3q$, then we need $l^{2}+1=0\pmod{4}$. Both obviously imply $l$ must be odd i.e., $l=2l^{\prime}+1$ for some integer $l^{\prime}$. But then it can easily be checked that $l^{2}+1=2\pmod{8}$ and thus $l^{2}+1\neq 0$ either modulo $4$ or $8$. So then we must have that $n=0,2q$. Clearly $n=0$ is simply $\mathbb{Z}_{2q}$ gauge theory, which is consistent with time reversal.
Suppose now that $n=2q$. Then our $K$-matrix is given by
@@
$W=\begin{pmatrix}1&0&0\\ 1&1&0\\ 0&0&1\end{pmatrix},\ W\in{\rm GL}(3,\mathbb{Z}),$ (94)
@@
Let $\mathcal{A}\in mG^{(\nu)}_{GSD}$ for the group $G$ in Eq. (4). Then it must be the case that $a^{2}$ is invariant under translation i.e., $T_{x}a^{2}=a^{2}=T_{y}a^{2}$. This then implies that $a^{2}$ is a boson. It will also further imply that the unique $\gamma$ discussed earlier is a boson.
@@
since $02\theta_{\gamma}=$ 0. Now we note that $a^{2q}=1$ so it must braid trivially with $a$, i.e., $0=\theta_{a,a^{2q}}=4q\theta_{a}$. Since $q$ is even this means that $(2q^{2}+4q)\theta_{a}=0$. Thus we have that $0=2\pi m/q$ so $m$ must be a multiple of $q$. The only anyons that $a$ can be taken to under translation are thus:
@@
In any case we see that $T_{x}a^{2}=(T_{x}a)(T_{x}a)=a^{2}$. The same logic will hold for $T_{y}$. Thus we see that $a^{2}$ must be invariant under translation.
Next we appeal to the logic of Appendix D.1. There we demonstrated that
@@
Let $\mathcal{A}\in mG^{(\nu)}_{\rm GSD}$ for the group $G$ in Eq. (4), where $q$ is even. Then $\mathcal{A}$ must be $\mathbb{Z}_{2q}$ gauge theory stacked with the fundamental fermion.